OpenStudy (anonymous):

Help walk me through please? I fan and medal! A modle rockeet is launched from a roof into a large dield. The path of the rocket can be modled by the equation y= -0.04x²+8.3x+4.3, where x is the horizontal distance in meters from the starting point on the roof and y is the hight, in meters, of the rocket obove the ground. How far horizontally from its starting point will the rocket land? Round your answer to the nearest 100th meter.

OpenStudy (anonymous):

okay i have read this now 4 times. the only thing i can think to do is to set this equal to zero, since it hits the ground at \(y=0\)

OpenStudy (anonymous):

that is, solve \[ -0.04x²+8.3x+4.3=0\] for \(x\)

OpenStudy (anonymous):

probably easiest to multiply by -100 and solve \[4x^2-830x-430=0\] then i guess the quadratic formula will give you the answer

OpenStudy (anonymous):

I got 0.52 m but i dont think thats right

OpenStudy (phi):

You should get 2 answers.

OpenStudy (anonymous):

no. the question was 'How far horizontally from its starting point will the rocket land'

OpenStudy (phi):

Let me be more clear. the quadratic equation -0.04x²+8.3x+4.3=0 has two solutions. One of them will be negative. One way to "get a feel" for the answer is to sketch the curve